Determine the quadratic function f whose graph is given.  
The vertex is
left parenthesis negative 3 comma negative 13 right parenthesis
(−3,−13) and the​ y-intercept is
negative 4.
−4.

Answer by stanbon(75887)

etermine the quadratic function f whose graph is given.  
The vertex is (−3,−13) and the​ y-intercept is −4.
--------
Form:: y = ax^2+bx+c
----
Since f(0) = -4, c = -4
-----
y = ax^2 + bx - 4
----
Since vertex is at -3, -3 = -b(2a)
Since Vertex is (-3,-13), -13 = a(-3)^2 + b(-3) - 4
------------
Solve for "a" and for "b"::
9a - 3b = -9
-b = -6a
-----
3a - b = -3
----
Therefore, 3a -6a = -3
-3a = -3
a = 1
Then b = 6a = 6
------
Ans: y = x^2 + 6x - 4
-----------------
Cheers,
Stan H.
-----------------